# Preliminaries on Pre-Hilbert Structures on Polynomial Spaces and Associated Laplacians

### v1

@@ 2.1. Polynomial spaces and pre-Hilbert structures
Let $\Omega\subset\mathbb{R}^{d}$ be a domain, and denote by $\mathcal{P}(\Omega)$ the space of (real or complex) polynomial functions restricted to $\Omega$. We consider a sesquilinear form
@@
which is assumed to be positive semi-definite. Such a form endows $\mathcal{P}(\Omega)$ with a pre-Hilbert structure. The completion (when Hausdorff) will be denoted by $H$.
@@
    $\langle p,qImgle_{L^{2}(\mu)}=\int_{\Omega}p(x)\overline{q(x)}\,d\mu(x),$
@@
  • Sobolev or fractional Sobolev inner products involving derivatives, extensively studied in [14];
  • discrete inner products arising from finite point sets and interpolation problems [1, 24].
Given such a pre-Hilbert structure, one may construct an orthogonal basis of $\mathcal{P}(\Omega)$ by applying the Gram–Schmidt procedure to the filtration by total degree.
2.2. Derivations and adjoint operators
Let $D$ denote a derivation on $\mathcal{P}(\Omega)$, typically the gradient operator $D=\nabla$. If $D$ is closable (and densely defined) in the Hilbert completion $H$, its adjoint $D^{*}$ is defined in the usual sense of unbounded operators on Hilbert spaces; see, e.g., [20, 8]. The operator
$\Delta:=D^{*}D$
is then a densely defined, positive, symmetric operator, which is self-adjoint under standard assumptions (e.g. via Friedrichs extension).
In the case where $\langle\cdot,\cdot\rangle$ is induced by a measure $\mu=\rho\,dx$, $\Delta$ coincides (up to sign conventions) with the weighted Laplacian
@@ 3. The Laplacian associated with a polynomial inner product
3.1. Closability of derivations and definition of $\Delta$
Let $\mathcal{P}(\Omega)$ be the space of (real or complex) polynomials restricted to a domain $\Omega\subset\mathbb{R}^{d}$, equipped with a positive definite inner product $\langle\cdot,\cdot\rangle$. Let $H$ be the Hilbert completion of $\mathcal{P}(\Omega)$. Fix a derivation $D$ on $\mathcal{P}(\Omega)$, typically $D=\nabla$ (or a finite family of derivations). Assume $D$ is closable in $H$ (this is automatic in many measure-induced or Sobolev settings, and is standard in the framework of closed forms; see [5, 18, 8, 17, 11]).
Remark 3.1.
In a more general setting, one can assume that $\langle\cdot,\cdot\rangle$ is only positive semi-definite, and that $\ker\langle\cdot,\cdot\rangle$ is invariant under $D$. Then $D$ descends to a well-defined derivation on $\mathcal{P}(\Omega)/\ker\langle\cdot,\cdot\rangle$. But this level of generality is not needed for our main examples. Therefore we choose to lead our main investigations for a positive definite inner product $\langle\cdot,\cdot\rangle$.
Denote by $\overline{D}$ its closure (still denoted $D$ for simplicity), and by $D^{*}$ its Hilbert adjoint. We define the associated (positive) Laplacian as the self-adjoint operator
(3.1) $\Delta:=D^{*}D,$
understood as the Friedrichs extension when $D^{*}D$ is initially defined only as a densely defined symmetric positive operator (standard; cf. [8, 18]).
@@ 3.3. Matrix representation in an orthonormal polynomial basis
(3.5) $\Delta_{m,n}:=\langle\Delta P_{n},P_{m}Imgle=\langle DP_{n},DP_{m}Imgle.$
@@ 3.4. Banded recurrences and banded Laplacians
Assume the operator matrix of $D$ in the orthonormal polynomial basis $(P_{n})$ is banded with lower bandwidth $r\geq 1$, i.e.
$\langle DP_{n},DP_{m}Imgle=0\quad\text{whenever}\quad|n-m|>r.$
Then the matrix $\Delta]=[\langle\Delta P_{n},P_{m}Imgle]_{m,n\geq 0}$ is banded with bandwidth at most $2r$, and is symmetric/Hermitian positive semidefinite.
@@
If $D$ is represented by a banded matrix $B$, then $\Delta=B^{*}B$. The product of a banded matrix with its adjoint has bandwidth at most the sum of bandwidths, hence at most $2r$. Hermitian positivity follows from $\Delta=B^{*}B$. ∎
@@ 3.5. Example: thin annuli and mode-by-mode block structure
(3.7) $\langle f\circ R_{\phi},g\circ R_{\phi}\rangle_{\varepsilon}=\langle f,gImgle_{\varepsilon},\qquad\forall f,g\in\mathcal{P}(A_{\varepsilon}),\ \forall\phi\in\mathbb{R}.$
@@
$\langle U_{\phi}f,U_{\phi}gImgle_{\varepsilon}=\langle f,gImgle_{\varepsilon},$
@@
$\langle D(U_{\phi}f),DgImgle_{\mathrm{vec}}=\langle U_{\phi}^{\mathrm{vec}}Df,DgImgle_{\mathrm{vec}}=\langle Df,(U_{\phi}^{\mathrm{vec}})^{-1}DgImgle_{\mathrm{vec}}=\langle Df,D(U_{-\phi}g)Imgle_{\mathrm{vec}}=\langle\Delta f,U_{-\phi}gImgle=\langle U_{\phi}\Delta f,gImgle.$
@@ 4.1. Weighted $L^{2}$ space and adjoint of the derivative
$\langle f,gImgle_{w}:=\int_{I}f(x)\,\overline{g(x)}\,w(x)\,dx.$
@@
(4.1) $\langle f^{\prime},gImgle_{w}=-\langle f,g^{\prime}+(w^{\prime}/w)gImgle_{w}+\big[f\,\overline{g}\,w\big]_{\partial I},$
@@ 4.2. Matrix representation in the orthonormal polynomial basis
(4.3) $(\Delta_{w})_{mn}:=\langle\Delta_{w}p_{n},p_{m}Imgle_{w}=\langle p_{n}^{\prime},p_{m}^{\prime}Imgle_{w}=\int_{I}p_{n}^{\prime}(x)\,\overline{p_{m}^{\prime}(x)}\,w(x)\,dx,$
@@
(4.4) $p_{n}^{\prime}(x)=\sum_{k=0}^{n-1}b_{n,k}\,p_{k}(x),\qquad b_{n,k}:=\langle p_{n}^{\prime},p_{k}Imgle_{w},$
@@ 5.1. Polynomial Hilbert geometries
Let $\Omega\subset\mathbb{R}^{d}$ be a domain and let $\mathcal{P}(\Omega)$ denote the space of polynomial functions restricted to $\Omega$.
@@
A polynomial Hilbert geometry on $\Omega$ is a triple
@@
  • $\langle\cdot,\cdot\rangle$ is a positive semi-definite sesquilinear form on $\mathcal{P}(\Omega)$;
  • $H$ denotes the Hilbert completion of $\mathcal{P}(\Omega)/\ker\langle\cdot,\cdot\rangle$;
  • $D:\mathcal{P}(\Omega)\to H^{m}$ is a derivation (typically $D=\nabla$), assumed to be closable in $H$.
@@
Two polynomial Hilbert geometries $\mathfrak{G}=(\mathcal{P},\langle\cdot,\cdot\rangle,D)$ and $\mathfrak{G}^{\prime}=(\mathcal{P},\langle\cdot,\cdot\rangle^{\prime},D)$ are said to be unitarily equivalent if there exists a unitary operator $U:H\to H^{\prime}$ such that:
@@
    $U$ maps $\mathcal{P}(\Omega)$ onto itself;

### v2

@@ 2.1. Polynomial spaces and pre-Hilbert structures
Let $\Omega\subset\mathbb{R}^{d}$ be a domain (i.e. an arbitrary subset, not necessarily open). We denote by
$\mathcal{P}(\Omega):=\{\,p|_{\Omega}:\ p\in\mathbb{C}[x_{1},\dots,x_{d}]\,\}$
the space of polynomial functions on $\Omega$. Let
@@
be a positive semidefinite sesquilinear form (conjugate-linear in the first variable). Set
$N:=\{p\in\mathcal{P}(\Omega):\langle p,p\rangle=0\},\qquad\mathcal{H}_{0}:=\mathcal{P}(\Omega)/N.$
Then $\langle\cdot,\cdot\rangle$ descends to a genuine Hermitian inner product on $\mathcal{H}_{0}$, and we denote by $H$ the Hilbert completion of $(\mathcal{H}_{0},\langle\cdot,\cdot\rangle)$. We will systematically identify $\mathcal{H}_{0}$ with its canonical dense image in $H$.
@@
    $\langle p,q\rangle_{L^{2}(\mu)}=\int_{\Omega}p(x)\,\overline{q(x)}\,d\mu(x),$
@@
  • Sobolev or fractional Sobolev inner products involving derivatives or nonlocal energies, see e.g. [14, 16];
  • discrete inner products arising from finite point sets and interpolation problems, [1, 24].
2.2. Gram–Schmidt orthogonalization and polynomial projectors
Fix an ordered family $(E_{n})_{n\geq 0}$ in $\mathcal{P}(\Omega)$ compatible with the degree filtration, for instance an enumeration of monomials with nondecreasing total degree. Assume that for each $N$ the family $(E_{0},\dots,E_{N})$ is linearly independent in $\mathcal{H}_{0}$. Define inductively the orthogonalized vectors $(\widetilde{p}_{n})_{n\geq 0}$ by
(2.1) $\widetilde{p}_{0}:=E_{0},\qquad\widetilde{p}_{n}:=E_{n}-\sum_{k=0}^{n-1}\frac{\langle E_{n},\widetilde{p}_{k}\rangle}{\langle\widetilde{p}_{k},\widetilde{p}_{k}\rangle}\,\widetilde{p}_{k},\quad n\geq 1,$
and the orthonormal polynomials $(p_{n})_{n\geq 0}$ by
(2.2) $p_{n}:=\frac{\widetilde{p}_{n}}{\|\widetilde{p}_{n}\|},\qquad\|\widetilde{p}_{n}\|^{2}=\langle\widetilde{p}_{n},\widetilde{p}_{n}\rangle.$
For each $N$, the subspace
$H_{\leq N}:=\mathrm{span}\{p_{0},\dots,p_{N}\}\subset H$
coincides with the image in $H$ of the polynomials of degree $\leq N$ (modulo $N$), and is finite-dimensional.
Gram matrices. Let $G_{N}\in\mathbb{C}^{(N+1)\times(N+1)}$ be the Gram matrix of $(E_{0},\dots,E_{N})$,
(2.3) $(G_{N})_{ij}:=\langle E_{j},E_{i}\rangle,\qquad 0\leq i,j\leq N.$
Then $G_{N}$ is Hermitian positive definite under the independence assumption above. Moreover, the coefficients in (2.1) can be written in matrix form: writing $E=(E_{0},\dots,E_{N})^{\top}$ and $\widetilde{p}=(\widetilde{p}_{0},\dots,\widetilde{p}_{N})^{\top}$, there exists a unique lower triangular matrix $L_{N}$ with positive diagonal such that
(2.4) $G_{N}=L_{N}L_{N}^{*},\qquad\widetilde{p}=L_{N}^{-1}E,$
and the normalized basis is $p=D_{N}^{-1}\widetilde{p}$ where $D_{N}=\mathrm{diag}(\|\widetilde{p}_{0}\|,\dots,\|\widetilde{p}_{N}\|)$.
Orthogonal projector. Let $P_{N}:H\to H_{\leq N}$ denote the orthogonal projector onto $H_{\leq N}$. In terms of the orthonormal basis $(p_{0},\dots,p_{N})$, one has for $f\in H$,
(2.5) $P_{N}f=\sum_{k=0}^{N}\langle f,p_{k}\rangle\,p_{k}.$
Equivalently, if we expand $P_{N}f=\sum_{j=0}^{N}c_{j}E_{j}$ in the (non-orthonormal) basis $(E_{0},\dots,E_{N})$, then the coefficient vector $c=(c_{0},\dots,c_{N})^{\top}$ is characterized by the normal equations
(2.6) $G_{N}\,c=b,\qquad b_{i}:=\langle f,E_{i}\rangle,\ 0\leq i\leq N,$
so $c=G_{N}^{-1}b$.
Finite-degree kernel (on compacts). For $x,y\in\Omega$, define the degree-$N$ kernel associated with $(p_{k})_{k=0}^{N}$ by
(2.7) $K_{N}(x,y):=\sum_{k=0}^{N}p_{k}(x)\,\overline{p_{k}(y)}.$
On a compact set $K\subset\Omega$, all polynomials are continuous, hence $K_{N}$ is continuous on $K\times K$. Moreover, $K_{N}$ is the matrix coefficient of $P_{N}$ evaluated on point-masses, in the sense made precise in Corollary 6.5 (kernel stability on compacts).
2.3. Derivations, closability, and associated Laplacians
Let $\widetilde{D}$ be a derivation on the polynomial algebra $\mathcal{P}(\Omega)$ (typically $\widetilde{D}$$\nabla$), taking values in $H^{m}$ (for some $m\geq 1$) and assume that
(2.8) $N\subset\ker\widetilde{D}$,
so that $\widetilde{D}$ induces a well-defined linear map $D_{0}:\mathcal{H}_{0}\to H^{m}$. Assume that $D_{0}$ is densely defined and closable. Denote by $D:=\overline{D_{0}}$ its closure and by $D^{*}$ its Hilbert adjoint. The associated Laplacian is the nonnegative self-adjoint operator
(2.9) $\Delta:=D^{*}D.$
(As $D$ is closed and densely defined, $D^{*}D$ is automatically self-adjoint and nonnegative.)
In the case where $\langle\cdot,\cdot\rangle$ is induced by a measure $\mu=\rho\,dx$ and $\widetilde{D}$$\nabla$, one recovers (up to sign conventions) the weighted Laplacian
@@ 2.3. Derivations, closability, and associated Laplacians
2.4. Resolvents, norm-resolvent distance, and finite-degree compressions
Let $\Delta$ be a nonnegative self-adjoint operator on $H$. Its resolvent $(1+\Delta)^{-1}$ is a bounded operator on $H$ with $\|(1+\Delta)^{-1}\|\leq 1$. The norm-resolvent topology is defined by convergence in operator norm of resolvents,
$\Delta_{n}\to\Delta\quad\Longleftrightarrow\quad\|(1+\Delta_{n})^{-1}-(1+\Delta)^{-1}\|_{\mathcal{B}(H)}\to 0,$
see [8, 20].
For a fixed degree $N$, we will use two related finite-dimensional objects:
  • the truncated operator on $H_{\leq N}$,
    $\Delta^{(N)}:=P_{N}\,\Delta\big|_{H_{\leq N}}:H_{\leq N}\to H_{\leq N};$
  • the compressed resolvent on $H_{\leq N}$,
    $R^{\langle N\rangle}:=P_{N}(1+\Delta)^{-1}P_{N}\in\mathcal{B}(H_{\leq N}).$
In general $R^{\langle N\rangle}\neq(I+\Delta^{(N)})^{-1}$, unless $H_{\leq N}$ is invariant under $\Delta$. The stability results of the paper are formulated in terms of the resolvent and its compressions, which are the quantities directly controlled by norm-resolvent estimates.
@@ 3.1. Degree growth and admissible derivations
3.2. Closability of derivations and definition of $\Delta$
Let $\Omega\subset\mathbb{R}^{d}$ be a domain and let
$\mathcal{P}(\Omega):=\{\,p|_{\Omega}:\ p\in\mathbb{C}[x_{1},\dots,x_{d}]\,\}$
be the space of polynomial functions on $\Omega$. Let $\langle\cdot,\cdot\rangle$ be a positive semidefinite sesquilinear form on $\mathcal{P}(\Omega)$ (conjugate-linear in the first variable), and set
$N:=\{p\in\mathcal{P}(\Omega):\langle p,p\rangle=0\},\qquad\mathcal{H}_{0}:=\mathcal{P}(\Omega)/N,$
so that $\langle\cdot,\cdot\rangle$ induces an inner product on $\mathcal{H}_{0}$. Denote by $H$ the Hilbert completion of $(\mathcal{H}_{0},\langle\cdot,\cdot\rangle)$.
Derivations and descent to the quotient. Let $\widetilde{D}$ be an algebraic derivation on $\mathcal{P}(\Omega)$, typically $\widetilde{D}$$\nabla$ (or a finite family of derivations), taking values in $H^{m}$. We assume that
(3.3) $N\subset\ker\widetilde{D}$,
so that $\widetilde{D}$ induces a well-defined linear operator
$D_{0}:\mathcal{H}_{0}\to H^{m},\qquad D_{0}([p])=\widetilde{D}(p).$
Remark 3.1 (Positive definite case).
If $\langle\cdot,\cdot\rangle$ is positive definite on $\mathcal{P}(\Omega)$, then $N=\{0\}$ and the quotient step is unnecessary: one simply has $\mathcal{H}_{0}=\mathcal{P}(\Omega)$. In this situation (3.3) is automatic.
Closability and definition of the Laplacian. Assume that $D_{0}$ is densely defined in $H$ (which holds in all our main examples, since $\mathcal{H}_{0}$ is dense in $H$) and that $D_{0}$ is closable. We denote by
$D:=\overline{D_{0}}:\mathrm{Dom}(D)\subset H\to H^{m}$
its closure, which is a closed densely defined operator, and by $D^{*}$ its Hilbert adjoint. We then define the associated (positive) Laplacian by
(3.4) $\Delta:=D^{*}D.$
Since $D$ is closed and densely defined, the operator $D^{*}D$ is automatically self-adjoint and nonnegative.
Remarks on examples. In measure-induced settings (e.g. $\langle f,g\rangle=\int f\overline{g}\,d\mu$) and Sobolev-type settings, closability of $D_{0}$ is standard and can be formulated at the level of closed quadratic forms; see, e.g., [5, 18, 8, 17, 11].
@@ 3.3. Matrix representation in an orthonormal polynomial basis
(3.6) $\Delta_{m,n}:=\langle\Delta P_{n},P_{m}\rangle=\langle DP_{n},DP_{m}\rangle.$
@@ 3.4. Banded recurrences and banded Laplacians
Let $(P_{n})_{n\geq 0}$ be an orthonormal polynomial basis of $H$ and let $B=(B_{k,n})_{k,n\geq 0}$ be the matrix of $D$ in this basis,
$DP_{n}=\sum_{k\geq 0}B_{k,n}\,P_{k},\qquad B_{k,n}:=\langle DP_{n},P_{k}\rangle.$
Assume that $B$ is banded with half-bandwidth $r\geq 1$, i.e.
$B_{k,n}=0\quad\text{whenever}\quad|k-n|>r.$
Then the matrix of $\Delta=D^{*}D$ in the same basis,
$\Delta_{m,n}:=\langle\Delta P_{n},P_{m}\rangle=\langle DP_{n},DP_{m}\rangle,$
is banded with half-bandwidth at most $2r$, i.e.
$\Delta_{m,n}=0\quad\text{whenever}\quad|m-n|>2r.$
Moreover $\Delta]$ is Hermitian and positive semidefinite.
@@
Since $(P_{n})$ is orthonormal, the coefficient array $B$ satisfies
$DP_{n}=\sum_{k\geq 0}B_{k,n}P_{k},\qquad\|DP_{n}\|^{2}=\sum_{k\geq 0}|B_{k,n}|^{2}.$
For $m,n\geq 0$ we compute, using orthonormality,
$\Delta_{m,n}=\langle DP_{n},DP_{m}\rangle$ $=\Big\langle\sum_{k\geq 0}B_{k,n}P_{k},\ \sum_{\ell\geq 0}B_{\ell,m}P_{\ell}\Big\rangle$
$=\sum_{k,\ell\geq 0}B_{k,n}\,\overline{B_{\ell,m}}\ \langle P_{k},P_{\ell}\rangle=\sum_{k\geq 0}B_{k,n}\,\overline{B_{k,m}}.$
This is exactly the matrix product identity
$\Delta]=B^{*}B,\qquad\Delta_{m,n}=\sum_{k\geq 0}\overline{B_{k,m}}\,B_{k,n}.$
Now assume |mn|>2r|m-n|>2r. If for some $k$ one has both $B_{k,n}\neq 0$ and $B_{k,m}\neq 0$, then by the bandedness hypothesis we must have $|k-n|\leq r$ and $|k-m|\leq r$, hence
$|m-n|\leq|m-k|+|k-n|\leq r+r=2r,$
a contradiction. Therefore for every $k$ at least one factor in $B_{k,n}\,\overline{B_{k,m}}$ is zero, and the sum vanishes: $\Delta_{m,n}=0$. This proves the half-bandwidth bound $2r$.
Finally, $\Delta]=B^{*}B$ implies that $\Delta]$ is Hermitian and positive semidefinite, since for any finitely supported vector $c=(c_{n})$,
$\sum_{m,n}\Delta_{m,n}c_{n}\overline{c_{m}}=\langle Bc,Bc\rangle_{\ell^{2}}\geq 0.$
@@ 3.5. Example: thin annuli and mode-by-mode block structure
(3.8) $\langle f\circ R_{\phi},g\circ R_{\phi}\rangle_{\varepsilon}=\langle f,g\rangle_{\varepsilon},\qquad\forall f,g\in\mathcal{P}(A_{\varepsilon}),\ \forall\phi\in\mathbb{R}.$
@@
$\langle U_{\phi}f,U_{\phi}g\rangle_{\varepsilon}=\langle f,g\rangle_{\varepsilon},$
@@
$\langle D(U_{\phi}f),Dg\rangle_{\mathrm{vec}}=\langle U_{\phi}^{\mathrm{vec}}Df,Dg\rangle_{\mathrm{vec}}=\langle Df,(U_{\phi}^{\mathrm{vec}})^{-1}Dg\rangle_{\mathrm{vec}}=\langle Df,D(U_{-\phi}g)\rangle_{\mathrm{vec}}=\langle\Delta f,U_{-\phi}g\rangle=\langle U_{\phi}\Delta f,g\rangle.$
@@ 4.1. Weighted $L^{2}$ space and adjoint of the derivative
$\langle f,g\rangle_{w}:=\int_{I}f(x)\,\overline{g(x)}\,w(x)\,dx.$
@@
(4.1) $\langle f^{\prime},g\rangle_{w}=-\langle f,g^{\prime}+(w^{\prime}/w)g\rangle_{w}+\big[f\,\overline{g}\,w\big]_{\partial I},$
@@ 4.2. Matrix representation in the orthonormal polynomial basis
(4.3) $(\Delta_{w})_{mn}:=\langle\Delta_{w}p_{n},p_{m}\rangle_{w}=\langle p_{n}^{\prime},p_{m}^{\prime}\rangle_{w}=\int_{I}p_{n}^{\prime}(x)\,\overline{p_{m}^{\prime}(x)}\,w(x)\,dx,$
@@
(4.4) $p_{n}^{\prime}(x)=\sum_{k=0}^{n-1}b_{n,k}\,p_{k}(x),\qquad b_{n,k}:=\langle p_{n}^{\prime},p_{k}\rangle_{w},$
@@ 5.1. Polynomial Hilbert geometries
Let $\Omega\subset\mathbb{R}^{d}$ be a domain and let $\mathcal{P}(\Omega)$ denote the space of polynomial functions on $\Omega$ (i.e. restrictions of ambient polynomials). Building on the explicitly treated examples, we can now introduce a unifying geometric framework for polynomial inner products and their associated operators. Some overlap with the preliminary section is therefore intentional: we prefer to restate the key constructions in a self-contained manner, in order to keep the general setting clear and easy to follow.
@@
Let $\Omega\subset\mathbb{R}^{d}$ and let $\mathcal{P}(\Omega)$ be the space of (restrictions to $\Omega$ of) complex polynomials. A polynomial Hilbert geometry on $\Omega$ is a triple
@@
  • $\langle\cdot,\cdot\rangle$ is a positive semidefinite sesquilinear form on $\mathcal{P}(\Omega)$ (conjugate-linear in the first variable);
  • $N:=\{p\in\mathcal{P}(\Omega):\langle p,p\rangle=0\}$ and $\mathcal{H}_{0}:=\mathcal{P}(\Omega)/N$; the form $\langle\cdot,\cdot\rangle$ descends to a Hermitian inner product on $\mathcal{H}_{0}$;
  • $H$ denotes the Hilbert completion of $(\mathcal{H}_{0},\langle\cdot,\cdot\rangle)$;
  • $\widetilde{D}:\mathcal{P}(\Omega)\to H^{m}$ is a derivation (typically $\widetilde{D}$$\nabla$), and we assume $N\subset\ker\widetilde{D}$ so that $\widetilde{D}$ induces a well-defined linear map $D_{0}:\mathcal{H}_{0}\to H^{m}$ by $D_{0}([p])=\widetilde{D}(p)$. We assume $D_{0}$ is densely defined and closable.
We denote by $D:=\overline{D_{0}}$ its closure and set $\Delta:=D^{*}D$.
@@
Remark 5.2 (On the notion of domain and the role of the quotient).
In this paper, a “domain” is simply a subset $\Omega\subset\mathbb{R}^{d}$, not necessarily open. We always interpret
$\mathcal{P}(\Omega):=\{\,p|_{\Omega}:p\in\mathbb{C}[x_{1},\dots,x_{d}]\,\}$
as the space of polynomial functions on $\Omega$ (restrictions of ambient polynomials). When $\Omega$ has empty interior (or more generally when the restriction map $p\mapsto p|_{\Omega}$ is not injective), distinct polynomials may induce the same function on $\Omega$. For instance, for $\Omega=S^{1}\subset\mathbb{R}^{2}$ one has the relation $x^{2}+y^{2}=1$ on $\Omega$, so the ideal $\langle x^{2}+y^{2}-1\rangle$ lies in the kernel of any $L^{2}(S^{1})$-type inner product. This explains the need to pass to the quotient $\mathcal{P}(\Omega)/\ker\langle\cdot,\cdot\rangle$.
@@
Two polynomial Hilbert geometries $\mathfrak{G}=(\mathcal{P}(\Omega),\langle\cdot,\cdot\rangle,D)$ and $\mathfrak{G}^{\prime}=(\mathcal{P}(\Omega),\langle\cdot,\cdot\rangle^{\prime},D^{\prime})$ are said to be unitarily equivalent if there exists a unitary operator $U:H\to H^{\prime}$ such that:
@@
    $U(\mathcal{H}_{0})=\mathcal{H}_{0}^{\prime}$ (identified as dense subspaces of $H$ and $H^{\prime}$);